# Memory Is Not the Bottleneck: Cost-Efficient Continual Learning via Weight Space Consolidation

### v3

@@ 4 Weight Space Consolidation
Based on insights from the previous section, we propose a cost-efficient CL method—Weight Space Consolidation—that leverages weight space operations (e.g., selective resets and running averages) to reconcile the trade-offs of the sufficient memory regime: high stability and low plasticity.
4.1 Rank-based parameter reset for improved plasticity
The core idea behind our method is that sufficient exemplar memory enables the model to start from a stable initialization (as discussed in Section 3), but remaining too close to this point can hinder plasticity. To address this, we introduce a rank-based reset technique that selectively reinitializes less important parameters based on their estimated contribution to learning.
Before training on the $t^{\text{th}}$ task, the model parameters—optimized on previous tasks $11{:}t{-}$ 1—are stored as $\theta_{\textrm{prev}}$ (see Line 6 in Algorithm 1). Training on the $t^{\text{th}}$ task proceeds using the loss for a task $\ell$ (Line 9), and after $n_{\textrm{warm}}$ warm-up epochs, we identify dormant parameters and gently reset them (Lines 10–11).
To rank parameter importance, we compute a moment-based metric $\mathcal{S}_{l}$ for each parameter element $l$, using the first and second exponential moments of its stochastic gradients ($\hat{m}_{l}$, $\hat{v}_{l}$), as in Adam [35]:
@@ 4.1 Rank-based parameter reset for improved plasticity
This score favors parameters that consistently receive strong gradients, while penalizing those with low or noisy updates. Specifically, a large $\hat{m}_{l}$ indicates that the parameter has received gradients in a consistent direction, while a large $\hat{v}_{l}$ implies that the parameter has experienced high gradient energy overall. By taking the product of the two, $\mathcal{S}_{l}$ becomes sensitive to both focused and sustained learning signals. Conversely, a low $\mathcal{S}_{l}$ suggests that the parameter has received weak or noisy gradients, indicating limited contribution to learning. We treat such parameters as dormant and reinitialize them to recover plasticity. This formulation provides a simple yet effective heuristic for identifying underutilized parts of the network based on gradient dynamics during warm-up.
In implementation, we retain only the top-$q\%$ (with $q{=}20$ by default following [68]), and reset the rest by softly blending them with $\theta_{\textrm{prev}}$ (Lines 12–13):
@@
1 Input: Model parameters $\theta$, training data $D_{1:t}$, memory buffer $\mathcal{M}$, average interval $j$, warming epoch $n_{\mathrm{warm}}$
@@ 11                   $\mathcal{I}_{\mathrm{reset}}\leftarrow\texttt{FindDormantParams}(\theta,\theta_{\mathrm{prev}})$
13                         Reset weights using eq. 5
14
15
@@ 20      $\theta\leftarrow\Theta$
Algorithm 1 Weight-space consolidation for cost-efficient CL
@@ 4.2 Weight averaging for improved stability
Using the reset model as a fresh starting point, we resume training $\theta$ for the remaining epochs. From this point on, we accumulate a running weight average $\Theta$ (see Lines 14–16) following the stochastic weight averaging (SWA) [28], which is known to promote convergence to flatter optima. The running average is updated every $j$ iterations after a warm-up phase of $n_{\textrm{warm}}$ steps:
@@
We find that this averaging is particularly effective under sufficient exemplar memory settings, where data diversity introduces significant gradient variance. After the warm-up phase, the model often oscillates around multiple distinct low-loss regions due to this variance. By averaging the weights across these regions, $\Theta$ converges to a flatter and more robust minimum that consolidates knowledge across both past and current tasks [28, 9].
Importantly, our approach differs from traditional model merging methods in continual learning, which often combine independently trained task-specific models to construct a final model [27]. In contrast, our method performs in-situ averaging during the training of a single model $\theta$, progressively updating $\Theta$ as a byproduct of the optimization trajectory. This eliminates the need to store and merge multiple per-task models, making our approach more memory-efficient and scalable to longer task sequences.
At the end of training on task $t$, we replace the model parameters with the averaged weights $\Theta$ (see Line 17), which then serve as a stable initialization for the next task, preserving knowledge while enabling further adaptation. Please refer to Section A.2 for further implementation details.
@@
Our method combines two simple yet effective weight-space operations to balance the stability–plasticity trade-off in the sufficient memory regime: (1) rank-based resets recover plasticity by reinitializing dormant parameters, and (2) weight averaging enhances stability by converging to flat, robust minima. Both are directly motivated by our analysis in Section 3 and introduce negligible overhead, requiring no storage of per-task models or additional backward passes.
@@ 5 Experiment
We aim to answer the following questions to assess the effectiveness, efficiency, and generality of our proposed method: (1) Does it outperform existing CL algorithms under sufficient memory? (2) Is it more cost-efficient than state-of-the-art algorithms like expansion- or merging-based approaches? (3) Can it generalize to other domains like continual learning for foundation models, e.g., continual instruction tuning for LLMs? (4) How do key design components—reset, averaging, importance metric—contribute to its effectiveness?
@@ 5.1 Experimental settings
We evaluate whether our method achieves strong performance and cost-efficiency under sufficient memory conditions across both vision and language domains.
Class-IL benchmarks.   We use two standard class-incremental learning benchmarks [47] via the PyCIL framework [78]: CIFAR-100, an image classification dataset with 100 classes split into 10 sequential tasks (10 classes each), and ImageNet-100, a 100-class subset of ImageNet also split into 10 tasks of 10 classes each. We compare our method to seven exemplar-based class-IL baselines: iCaRL [54], BiC [67], WA [76], DER [69], FOSTER [61], MEMO [80], and Replay, a naive baseline that finetunes using current data and stored exemplars.
LLM continual instruction tuning.   We also evaluate on TRACE [64], a continual instruction tuning benchmark for LLMs across eight domains, including science [43], policy [57], meeting summarization [24], multilingual classification [77, 20], code generation [44], and math reasoning [50]. We compare our method against Replay and four model-merging baselines, motivated by recent findings that such methods can be effectively applied to continual learning for foundation models [56, 15]: Model Soups [66], SLERP [30], MagMax [45], and Task Arithmetic [27].
Architectures and protocol.   For image classification, we use ResNet-32 [23] on CIFAR-100 and ResNet-18 [23] on ImageNet-100 [23]. For LLM experiments, we use an instruction-tuned version of LLaMA-3.2 [21]. All class-IL results are averaged over five seeds and reported as average class-IL accuracy across all tasks after the final step [47]. For TRACE, final scores are reported after completing all sequential tasks. We select hyperparameters following the realistic CL evaluation protocol proposed by [7]. Further details on experimental settings are in Section A.5.
@@ 5.2 Experimental results
Class-IL results.   We report class-incremental learning results in Table 1 and Figures 1, 11. Note that DER, FOSTER, and MEMO are expansion-based methods that increase model size over time during training. To ensure a fair comparison, we follow the evaluation protocol of [7] and reuse the best hyperparameters found on CIFAR-100 when evaluating on ImageNet-100. From these results, we make three key observations: First, as shown in Table 1, under the conventional constrained memory setting (e.g., 4% memory), existing class-IL methods outperform Replay. However, as the memory size increases, the performance gap narrows substantially. With 20% memory, most methods perform similarly to Replay, suggesting diminishing returns of algorithmic complexity in the sufficient memory regime. Second, when training cost (i.e., training time) is taken into account, Figures 1 and 11 (see Section A.4) show that expansion-based methods become highly inefficient. For instance, while FOSTER maintains strong accuracy even under sufficient memory, its training time is 4–5$\times$ higher than that of Replay. Third, our proposed method—Weight Space Consolidation—demonstrates both strong performance and high efficiency. Table 1 shows that it consistently outperforms existing baselines (except for expansion-based methods) under sufficient memory (i.e., over 20% memory). Meanwhile, note that its training cost remains comparable to that of Replay, as shown in Figures 1 and 11. Together, these results validate that our method effectively mitigates the plasticity–stability trade-off in class-IL using sufficient memory with minimal GPU computation overhead.
Table 1: Average class-IL accuracy (%) on CIFAR-100 and ImageNet-100 with varying exemplar-memory sizes. We report experimental results with varying memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 400/600 exemplars per class (storing $80\%$ of the previous dataset in CIFAR-100 and nearly half in ImageNet-100). Bold highlights the best non-expansion method.
@@
Continual instruction tuning results.   We evaluate our method in a more practical setting of continual learning for foundation models, where exemplar memory is sufficient. To this end, we conduct continual instruction tuning on the TRACE benchmark, following its standard 8-stage setup across diverse domains. As strong baselines, we compare with recent model merging approaches that have been actively explored for continual learning with foundation models [56, 15]. Among them, methods such as Task Arithmetic, MagMax, and Model Soup (i.e., Avg. (Offline)) require storing all $T$ task-specific models or representations and merging them post hoc. In contrast, methods like Ours, SLERP, and Avg. (Online) operate in a single model trajectory and do not require saving all intermediate checkpoints. We denote $|W|$ as the standard vram usage of a single model. To ensure a fair evaluation, we follow the protocol proposed by [7]: all hyperparameters are tuned only under the 20% memory setting, and the same configuration is reused across other memory sizes without additional tuning. Figure 3 shows that our method consistently outperforms the baselines when the exemplar memory exceeds 20%. In particular, we observe that our method achieves 2–9% higher accuracy across all memory sizes compared to offline merging approaches like Task Arithmetic and MagMax, which require a relative vram usage of $|W|*T$ for saving all previous task models in the GPU. Notably, as summarized in Appendix Table 7, these offline merging methods struggle to integrate knowledge from diverse tasks [15] and tend to underperform under sufficient memory conditions. Conversely, our method avoids these issues by operating directly in weight space during training and achieves both high performance and GPU efficiency (relative vram usage: $|W|*2$).
More experimental results or detailed numerical experimental results are introduced in Section A.4.
@@ 5.3 Ablation study
Table 2: Ablation study on CIFAR-100
@@
We conduct ablation studies to validate the effectiveness of each component of our method. Our approach consists of two key weight-space operations: (1) rank-based parameter reset for plasticity recovery, and (2) weight averaging for stability.
Contribution of each component.   Table 2 summarizes the performance of our method when either reset or averaging is ablated. The baseline Replay uses neither. We find that weight reset alone (w/o avg.) yields limited gains over Replay, while averaging without reset (w/o reset) improves stability but fails to fully adapt to new tasks. Only when both are combined do we observe substantial performance gains. This highlights that the two operations are complementary and that weight reset primarily serves to enable effective averaging.
Sparsity analysis of reset.   Next, we examine how much of the model contributes to learning new tasks. In Figure 5, we vary the retain rate $Q$ in the reset step. We find that resetting up to 80% of parameters yields minimal degradation in accuracy, suggesting that only a small subset of weights are actively involved in continual learning. This aligns with findings in sparse training and pruning [19, 11]. We observe that weight averaging helps retain performance under extreme resets (e.g., $Q{=}10\%$), suggesting that it encourages robust, sparse models.
Effectiveness of the importance metric. Table 4 compares alternative parameter importance metrics used for the reset step. Our moment-based metric (Eq. 4) performs on par with expensive Hessian-based scores, while being significantly more efficient. Simple metrics like parameter drift also perform well, echoing prior work in pruning [81, 42]. In contrast, using only the first or second moment leads to poor selection, confirming the necessity of combining both for robustness.
@@
Finally, we investigate two design choices in the reset operation: the reset rule (how to reset) and the reset frequency (when to reset). As shown in Table 4, our soft reset method (a weighted blend with $\theta_{\text{prev}}$) consistently outperforms random reinitialization and hard reversion. This advantage becomes more pronounced with larger memory. Regarding frequency, performing a reset only once after $n_{\text{warm}}$ works well in most settings. However, in extremely constrained memory regimes, applying multiple resets yields further gains—consistent with findings in sparse reinitialization [19]. We provide further analysis in Section A.4 on how tuning the reset frequency and retain rate $Q$ affects performance.
Table 3: Study of Parameter-Importance Metric on Average class-IL accuracy (%) in CIFAR-100.
Memory Size Metric Cost 20 200 500 Param. Drift ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\Downarrow}}$ 51.93 73.02 77.90 Fisher-based ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\Uparrow}}$ 52.35 73.49 77.57 Hessian-based ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\Uparrow}}$ 52.81 73.63 78.16 First Moment ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\Downarrow}}$ 50.10 71.38 75.05 Second Moment ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\Downarrow}}$ 47.29 70.44 73.89 Ours (eq. 4) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\Downarrow}}$ 52.00 73.57 78.25
Table 4: Study of reset strategies on Average class-IL accuracy (%) in CIFAR-100.
Memory Size Strategy 20 200 500 Reset Method Random 45.30 71.56 75.11 Revert 50.93 71.85 75.78 Mix (Ours) 52.00 73.57 78.25 Reset Frequency Every Iter. 47.38 63.11 70.32 Every Epoch 53.36 72.73 76.57 Once (Ours) 52.00 73.57 78.25

### v5

@@ 4 Weight Space Consolidation
Based on insights from the previous section, we propose a cost-efficient CL method—Weight Space Consolidation—that leverages weight space operations (e.g., selective resets and running averages) to reconcile the trade-offs of the sufficient exemplar memory regime: high stability and low plasticity.
4.1 Ranking-based parameter reset for improved plasticity
The core idea behind our method is that sufficient exemplar memory enables the model to start from a stable initialization (as discussed in Section˜3), but remaining too close to this point can hinder plasticity. To address this, we introduce a ranking-based reset technique that selectively reinitializes less important parameters based on their estimated contribution to learning.
Before training on the $t^{\text{th}}$ task, the model parameters—optimized on previous tasks 1:t11{:}t{-}1—are stored as $\theta_{\textrm{prev}}$ (Line 6 in Algorithm˜1). Training on the $t^{\text{th}}$ task proceeds using the task loss $\ell$ (Line 9), and after $n_{\textrm{warm}}$ warm-up epochs, we identify dormant parameters and gently reset them (Lines 10–11).
To rank parameter importance, we compute a moment-based metric $\mathcal{S}_{l}$ for each parameter element $l$, using the first and second exponential moments of its stochastic gradients ($\hat{m}_{l}$, $\hat{v}_{l}$), as in Adam (Kingma and Ba, 2017):
@@ 4.1 Ranking-based parameter reset for improved plasticity
This score favors parameters that consistently receive strong gradients, while penalizing those with low or noisy updates. Specifically, a large $\hat{m}_{l}$ indicates that the parameter has received gradients in a consistent direction, while a large $\hat{v}_{l}$ implies that the parameter has experienced high gradient energy overall. By taking the product of the two, $\mathcal{S}_{l}$ becomes sensitive to both focused and sustained learning signals. Conversely, a low $\mathcal{S}_{l}$ suggests that the parameter has received weak or noisy gradients, indicating limited contribution to learning. We treat such parameters as dormant and reinitialize them to recover plasticity. This formulation provides a simple yet effective heuristic for identifying underutilized parts of the network based on gradient dynamics during warm-up. In Section˜5, we compare existing methods that use weight resets for plasticity recovery (see Table˜5).
In implementation, we retain only the top-$Q\%$ of the parameter element $l$ (with Q=20Q{=}20 by default following (Yadav et al., 2024)), and reset the rest by softly blending them with $\theta_{\textrm{prev}}$ (Lines 12–13):
@@
1Input: Model parameters $\theta$, training data $D_{1:t}$, memory buffer $\mathcal{M}$, average interval $j$, warming epoch $n_{\mathrm{warm}}$
@@ 12       $\mathcal{I}_{\mathrm{reset}}\leftarrow\texttt{FindDormantParams}(\theta,\theta_{\mathrm{prev}})$
14            Reset weights using eq.˜5
15         end for
16
17      end if
18
@@ 25 $\theta\leftarrow\Theta$
26 end for
Algorithm 1 Weight Space Consolidation for cost-efficient CL
@@ 4.2 Weight averaging for improved stability
Using the reset model as a fresh starting point, we resume training $\theta$ for the remaining epochs. From this point on, we accumulate a running weight average $\Theta$ (see Lines 14–16) following the stochastic weight averaging (SWA) (Izmailov et al., 2018), which is known to promote convergence to flatter optima. The running average is updated every $j$ iterations after a warm-up phase of $n_{\textrm{warm}}$ steps:
@@
We find that this averaging is particularly effective under sufficient exemplar memory settings, where data diversity introduces significant gradient variance. After the warm-up phase, the model often oscillates around multiple distinct low-loss regions due to this variance. By averaging weights across these regions, $\Theta$ converges to a flatter and more robust minimum that consolidates knowledge across both past and current tasks (Izmailov et al., 2018; Cha et al., 2020).
Importantly, our approach differs from traditional model merging methods in CL, which often combine independently trained task-specific models to construct a final model (Ilharco et al., 2022). In contrast, our method performs in-situ averaging during the training of a single model $\theta$, progressively updating $\Theta$ as a byproduct of the optimization trajectory. This eliminates the need to store and merge multiple per-task models, improving our method’s cost-efficiency and scalability to longer task sequences.
At the end of training on task $t$, we replace the model parameters with the averaged weights $\Theta$ (see Line 17), which then serve as a stable initialization for the next task, preserving knowledge while enabling further adaptation. Please refer to Section˜A.2 for further implementation details.
@@
Our method combines two simple yet effective weight-space operations to balance the stability–plasticity trade-off in the sufficient exemplar memory regime: (1) ranking-based resets recover plasticity by reinitializing dormant parameters, and (2) weight averaging enhances stability by converging to flat, robust minima. Both are directly motivated by our analysis in Section 3 and introduce negligible overhead, requiring no storage of per-task models or additional backward passes.
@@ 5.1 Experimental settings
We evaluate our method’s performance and cost-efficiency under various exemplar memory settings.
Class-IL benchmarks. We use two standard class-incremental learning benchmarks (Masana et al., 2022) via the PyCIL framework (Zhou et al., 2023a): CIFAR-100, an image classification dataset with 100 classes split into 10 sequential tasks (10 classes each), and ImageNet-100, a 100-class subset of ImageNet also split into 10 tasks of 10 classes each. We compare our method to seven exemplar-based class-IL baselines: iCaRL (Rebuffi et al., 2017), BiC (Wu et al., 2019), WA (Zhao et al., 2020), DER (Yan et al., 2021), FOSTER (Wang et al., 2022), MEMO (Zhou et al., 2023b), and Replay, a naive baseline that finetunes using current data and stored exemplars.
LLM continual instruction tuning. We also evaluate on TRACE (Wang et al., 2023), a continual instruction tuning benchmark for LLMs across 8 domains.We compare our method against Replay and 4 model-merging baselines, following recent findings that such methods can be effectively applied to CL (Roth et al., 2024; Dziadzio et al., 2024): Model Soups (Wortsman et al., 2022), SLERP (Jang et al., 2024), MagMax (Marczak et al., 2025), and Task Arithmetic (Ilharco et al., 2022).
Architectures and protocol. For image classification, we use ResNet-32 (He et al., 2016) on CIFAR-100 and ResNet-18 (He et al., 2016) on ImageNet-100 (He et al., 2016). For LLM experiments, we use an instruction-tuned version of LLaMA-3.2 (Grattafiori et al., 2024). All class-IL results are averaged over five seeds and reported as average class-IL accuracy across all tasks after the final step (Masana et al., 2022). For TRACE, final scores are reported after completing all sequential tasks. We select hyperparameters following the realistic CL evaluation protocol proposed by (Cha and Cho, 2025). Further details on experimental settings are in Section˜A.5.
@@ 5.2 Experimental results
Table 2: Average class‑IL accuracy (%) on CIFAR‑100 and ImageNet‑100 with varying exemplar‑memory sizes. We report experimental results with varying memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 400/600 exemplars per class (storing $80\%$ of the previous dataset in CIFAR-100 and nearly half in ImageNet-100). Bold highlights the best non‑expansion method. We report the standard error across 5 runs.
[…]
@@
Continual instruction tuning results. We evaluate our method in a more practical setting of continual learning for foundation models, where exemplar memory is abundant. To this end, we conduct continual instruction tuning on the TRACE benchmark, following its standard 8-stage setup across diverse domains. As strong baselines, we compare with recent model merging approaches that have been actively explored for continual learning with foundation models (Roth et al., 2024; Dziadzio et al., 2024). Among them, methods such as Task Arithmetic, MagMax, and Model Soup (i.e., Avg. (Offline)) require storing all $T$ task-specific models or representations and merging them post hoc. In contrast, methods like Ours, SLERP, and Avg.(Online) operate in a single model trajectory and do not require saving all intermediate checkpoints. We denote $|W|$ as the standard vram usage of a single model. To ensure a fair evaluation, we follow the protocol proposed by (Cha and Cho, 2025): all hyperparameters are tuned only under the 20% memory setting, and the same configuration is reused across other memory sizes without additional tuning. Figure˜3 shows that our method consistently outperforms the baselines when the exemplar memory exceeds 20% (Also see Figure˜10). In particular, we observe that our method achieves 2–9% higher accuracy across all memory sizes compared to offline merging approaches like Task Arithmetic and MagMax, which require a relative vram usage of |W|T|W|*T for saving all previous task models in the GPU. Notably, as summarized in Appendix Table˜9, these offline merging methods struggle to integrate knowledge from diverse tasks (Dziadzio et al., 2024) and tend to underperform under abundant exemplar memory conditions. Conversely, our method avoids these issues by operating directly in weight space during training and achieves both high performance and GPU efficiency (relative vram usage: |W|2|W|*2). More experimental results or detailed numerical experimental results are introduced in Section˜A.4.
@@ 5.3 Ablation study
Table 3: Ablation study on CIFAR‑100
@@
We conduct ablation studies to validate the effectiveness of two key components of our method. (1) ranking-based parameter reset for plasticity recovery, and (2) weight averaging for stability.
Contribution of each component. Table 3 summarizes the performance of our method when either reset or averaging is ablated. The baseline Replay uses neither. We find that weight reset alone (w/o avg.) yields limited gains over Replay, while averaging without reset (w/o reset) improves stability but fails to fully adapt to new tasks. Only when both are combined do we observe substantial gains, highlighting that the two operations are complementary and that weight reset primarily serves to enable effective averaging.
Table 4: Study of Parameter‐Importance Metric on Average class-IL accuracy (%) in CIFAR-100
[…]
Effectiveness of the importance metric. Table 4 compares alternative parameter importance metrics used for reset. Our moment-based metric (Eq. 4) performs on par with expensive Hessian-based scores, while being significantly more efficient. Simple metrics like parameter drift also perform well, echoing prior work in pruning (Zhu and Gupta, 2017; Liu et al., 2018). In contrast, using only the first or second moment leads to poor selection, confirming the necessity of combining both for robustness.
Table 5: Study of reset strategies on Average class-IL accuracy (%) in CIFAR-100
Memory Size Strategy 20 200 500 Reset Method Random 45.30 71.56 75.11 Revert 50.93 71.85 75.78 Shrink&Perturb 3 48.26 70.32 76.05 Continual Backprop. 16 47.15 70.49 76.52 Ours (w/o Avg.) 48.73 70.81 76.47 Ours 52.00 73.57 78.25 Reset Frequency Every Iter. 47.38 63.11 70.32 Every Epoch 53.36 72.73 76.57 Once (Ours) 52.00 73.57 78.25
@@
Finally, we investigate two design choices in the reset operation: the reset rule (how to reset) and the reset frequency (when to reset). As shown in Table 5, our soft reset method (a weighted blend with $\theta_{\text{prev}}$) consistently outperforms random reinitialization and hard reversion. This advantage becomes more pronounced with larger memory. We also compare with existing methods (e.g., Shrink & Perturb (Ash and Adams, 2020) and Continual Backprop. (Dohare et al., 2024)), where our method outperforms competitors. Regarding frequency, performing a reset only once after $n_{\text{warm}}$ works well in most settings. However, in constrained memory regimes, applying multiple resets yields further gains—consistent with findings in sparse reinitialization (Frankle and Carbin, 2019). We provide further analysis in Section˜A.4 on how tuning the reset frequency and retain rate $Q$ affects performance.
Table 6: Comparison of Full Retraining from scratch and Continual Learning under a standard CL setting. Average class-IL accuracy (%) on CIFAR-100 is reported.
[…]
Continual Learning vs. Full Retraining.
Prior work (Dohare et al., 2024) reports that with full exemplar memory, retraining from scratch (i.e., joint-training) can surpass continual learning (CL). In Section˜3.2, we revisit this observation theoretically and argue that, under full memory, per-task resets (reinitializing weights before each task) recover plasticity while preventing forgetting, yielding strong performance. However, in the more realistic sufficient memory regime (e.g., 20–40%), retraining from scratch degrades markedly as forgetting is insufficiently addressed. In these settings, algorithms that leverage prior models (e.g., ours) offer clear advantages, as seen in Table˜6. Finally, while maximal replay might appear ideal, it is often impractical at scale (e.g., LLMs). Our approach provides a cost-effective alternative that balances performance with efficiency